OpenStudy (anonymous):

Which shows solutions to x + 4y = 9 if x and y must be whole numbers. A. {(0, 9), (1, 4), (2, 1)} B. {(2, 1), (1, 6), (0, 9)} C. {(1, 2), (5, 1), (9,0)} D. {(0, 9), (1, 5), (2, 1)}

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

x+4y=9 So x=9-4y 1=9-4*2 5=9-4*1 9=9-4*0 So you are right!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome!